OpenStudy (anonymous):

What is the momentum of a proton (in GeV/c) with KE=800MeV?

OpenStudy (anonymous):

\[K =.8GeV\] \[ K=\frac{p^2}{2m_p}\] \[p=\sqrt{2Km_p}\] \[m_p=.938GeV/c^2\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

It's relativistic, so classical equations don't apply.

OpenStudy (anonymous):

I've been banging my head against this for an hour, I'd be very grateful if you actually helped me solve it.

OpenStudy (anonymous):

I'm pretty sure this equation is what's used to solve it: E^2=p^2c^2+m^2c^4

OpenStudy (anonymous):

E = total energy, which is kinetic energy + rest energy, which is 1739.57 MeV. But when I plug that into the equation and solve, I end up with something that's off by many orders of magnitude.

OpenStudy (anonymous):

The answer should be in GeV, and I'm ending with something that's closer to a single eV. I'm beginning to hate physics.

OpenStudy (anonymous):

hmmm, lemme see. 1 sec.... And physics LOVES you! It's just a coy mistress ^_^

OpenStudy (anonymous):

I don't like coy mistresses... >.>

OpenStudy (anonymous):

I like my mistresses straightforward and honest!

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

does 1.46 GeV/c sound right?

OpenStudy (anonymous):

That was right! How did you get it?

OpenStudy (anonymous):

She just needs to know that you understand her quirks ~_^ You had everything right, you just divided by the numerical value of c \[E^2 = p^2c^2+m^2c^4\] \[(1.739 GeV)^2=p^2c^2+(.938GeV/c^2)^2c^4\] \[(1.739 GeV)^2=p^2c^2+(.938GeV/c^2)^2c^4\] \[(1.739 GeV)^2=p^2c^2+(.938GeV)^2\] \[(1.739 GeV)^2-(.938GeV)^2=p^2c^2\] \[p=\sqrt{\frac{(1.739 GeV)^2-(.938GeV)^2}{c^2}}=\sqrt{(1.739 )^2-(.938)^2}\frac{GeV}{c}\] The c is actually a part of the units!

OpenStudy (anonymous):

(for that last step I just pulled the units out of the square root)

OpenStudy (anonymous):

Where did the c^4 go in the third step?

OpenStudy (anonymous):

Canceled with the c^2^2 I presume, but why didn't the exponent distribute to the .938 as well?

OpenStudy (anonymous):

Yeah, the c^4's canceled out. What do you mean?

OpenStudy (anonymous):

\[(.938 \frac{GeV}{c^2})^2c^4=(.938^2 \frac{GeV^2}{c^4})c^4 =(.938^2 \frac{GeV^2c^4}{c^4})=(.938^2 GeV^2) = (.938GeV)^2\]

OpenStudy (anonymous):

~last term \[=(.938 GeV)^2\]

OpenStudy (anonymous):

So why doesn't that become .938^2GeV?

OpenStudy (anonymous):

The GeV is also squared - there's nothing to cancel it out

OpenStudy (anonymous):

i.e. .879844GeV?

OpenStudy (anonymous):

if you wrote it like that, it would be \[ .8798 \ GeV^2\]

OpenStudy (anonymous):

just like if you had \[(10 m)^2 = (100m^2)\]

OpenStudy (anonymous):

Okay, I see, thanks a lot.

OpenStudy (anonymous):

^_^ welcome! Always glad to act as a relationship counselor ^_^

OpenStudy (anonymous):

Could you have solved this by leaving the mass in kg?

OpenStudy (anonymous):

You could go through all the steps, but to get units of GeV would would have at some point had to convert kg into GeV/c^2 So to do it that way, just leave all of the c's as the letter c, then your final expression would look like \[p=\sqrt{\frac{(1.739GeV)^2-(m_pc^2)^2}{c^2}}\] then \[1.782662×10^{−36} kg = 1 eV/c^2\] Conversely, if you do everything in terms of meters, seconds, and kilograms you have to convert MeV into Joules in the beginning (then your numbers are crazy tiny though)

OpenStudy (anonymous):

Are you still here?

OpenStudy (anonymous):

hi, yah. The site kicked me off for a minute.

OpenStudy (anonymous):

Me too, I guess they just did an update. Thanks for your help. I'm gonna close this one and ask another question.

OpenStudy (anonymous):

Cool cool, and again, welcome ^_^